A particle is moving in positive x-direction with initial velocity of 10 m/s and uniform retardation such that it reaches the initial position after 10s.The distance transversed by the particle in 6s is ?

Grade:10

1 Answers

Eshan
5 years ago
Dear student,

Froms=ut+\dfrac{1}{2}at^2=0\implies t=-\dfrac{2u}{a}
\implies a=-2m/s^2

Hences(t=6s)=ut+\dfrac{1}{2}at^2=60-36=24m